M olar mass of ethanol (C 2 H 5 OH ) is : 34.2 38.9 46.1 45.1 62.1 2

Molar mass of ethanol (C2H5OH)  is :

  1.  34.2 
  2.  38.9
  3.  46.1
  4.  45.1
  5.  62.1

2


 

Grade:12th pass

2 Answers

Vikas TU
3 years ago
Dear st6udent 
Molar mass of C2H5OH = 2(12) + 6(1)+16 = 46.1 g/mol 
 
Hope this will help 
Good Luck 
Cheers 
Vikas Amritiya
3 years ago
Dear student

Molar mass of ethanol (C2H5OH=2(12)+6(1)+16.1=46.1 g/mol

 
Hence the correct option is C.
Best wishes!